# Graph Network Simulator MGN MeshGraphNet

### v1

@@ Appendix E Results
Figure 9 visualize the displacement rollout RMSE over the inference time of on ImpactPlate and DeformingPlateROBIN is compared to different inference modes and trained variants with fewer diffusion steps.
@@
Figure 9: RMSE rollout error and inference time of ROBIN, different inference modes of ROBIN, and variants with fewer denoising steps $K$ on a) ImpactPlate and b) DeformingPlateROBIN is most accurate on ImpactPlate and on par with the slower variants ROBIN - $m5$ and ROBIN - $m20$. The one step variant ROBIN - OS achieves the largest speed-up on DeforminPlate, but also loses the most accuracy there. Reducing the diffusion steps to $10$ (ROBIN - $K10$) and $5$ (ROBIN - $K5$) denoising steps, respectively, increases speed while decreases the accuracy.
@@
Figure 10 and Figure 11 visualize the rollout displacement and von Mises stress prediction of ROBINHCMT and MGN on ImpactPlate and DeformingPlate, respectively.
@@
Figure 10: Comparison of the rollout deformation prediction and von Mises stress prediction (color) on ImpactPlate to the ground truth of the FEMROBIN most accurately resolves the deformation at the contact surface and the resulting stress. The deformation prediction of HCMT comes close to the FEM prediction while the stress is underestimated. MGNs predicts accurately the global modes but exhibits local disturbances.
@@
Figure 11: Rollout deformation and von Mises stress prediction (color) on DeformingPlate of ROBIN, the baselines and the FEM. All models accurately reproduce the part deformation. HCMT overestimates slightly the stress at the thin wall between the hole and the boundary.

### v2

@@ Appendix E Results
AMG-based mesh coarsening.
Root-node AMG coarsening [53] preserves mesh geometry and connectivity (cf. Figure˜10). The coarse mesh remains well-aligned with thin geometrical features, and smoothed transfer operators yield wider, algebraically informed receptive fields compared to bi-stride pooling. This fidelity is crucial for predicting geometrically nonlinear deformations.
Figure 10: Comparison of AMG-based mesh coarsening to BSDL. a) the original, fine mesh. b) the mesh after two BSDL coarsening steps. c) the mesh after one AMG coarsening step. This mesh has approximately as many nodes as the mesh in b). d) up- and downsampling edges from level 0 to 11 (top) and level $1$ to 22 (bottom). Bright blue indicates coarse nodes of level 11 (top) or 22 (bottom), respectively.
Inference speed.
Each point in Figure˜11 corresponds to a rollout setting of ROBIN on ImpactPlate and DeformingPlate. As for BendingBeam in Section˜5, decreasing the truncation step reduces wall time while a truncation step of $k_{\text{tr}}{=}2$ is already sufficient to obtain a higher accuracy than all baselines across all datasets (cf. Table˜3). The ROBIN default $(1/20)$ is significantly faster than conventional diffusion inference and is even slightly more accurate on ImpactPlate, which we attribute to anchoring low-frequency components and reducing drift.
@@
Figure 11: RMSE rollout error and inference time of different inference variants of ROBIN on a) ImpactPlate and b) DeformingPlateROBIN, i.e., the variant (1/20), is most accurate on ImpactPlate and on par on DeformingPlate with the slower variants (5/20) and conventional inference (20/20). Reducing the truncation step $k_{\text{tr}}$ increases speed while decreasing accuracy. The one step variant (1/1) achieves the largest speed-up on DeformingPlate, but also loses the most accuracy there.
Baselines.
@@
Figure˜12 and Figure˜13 visualize the rollout displacement and von Mises stress prediction of ROBINHCMTMGN, and BSMS on ImpactPlate and DeformingPlate, respectively.
@@
Figure 12: Comparison of the rollout deformation prediction and von Mises stress prediction (color, yellow is high) on ImpactPlate to the ground truth of the FEMROBIN most accurately resolves the deformation at the contact surface and the resulting stress. The deformation prediction of HCMT and BSMS are close to the FEM prediction, though stress is underestimated. MGN predicts accurately the global modes but exhibits local disturbances.
@@
Figure 13: Rollout deformation and von Mises stress prediction (color, yellow is high) on DeformingPlate of ROBIN, the baselines and the FEM. All models accurately reproduce the part deformation. HCMT slightly overestimates and BSMS slightly underestimates the stress at the thin wall between the hole and the boundary.
Diffusion truncation. The rollouts in Figure˜14 across truncation steps $k_{\text{tr}}$ illustrate coarse-to-fine frequency behavior of ROBIN. Early steps capture the global, low-frequency shape, whereas additional steps sharpen high-frequency details, e.g., high stresses in thin geometrical features. Even with early truncation, ROBIN maintains coherent global modes. Longer schedules primarily refine local features and minimize the accumulation of high-frequency errors, while retaining the global deformation pattern.
Figure˜15 visualize the rollout displacement RMSE and displacement gradient RMSE of different ROBIN variants on DeformingPlate and ImpactPlate. As for BendingBeam, early denoising steps are critical for reducing global displacement error. Later diffusion steps focus on high-frequency solution components. On ImpactPlate, the gradient RMSE shows a modest increase while ROBIN attains the lowest displacement RMSE overall. We hypothesize that partially denoising states (small $m$) stabilizes low-frequency components and reduces drift. In contrast, fully denoising (larger $m$) suppresses short-term, high-frequency error accumulation. However, this trade-off does not affect other datasets with significantly longer rollouts.
Figure 14: Effect of diffusion truncation on rollouts. Figures a), b), and c) show predicted deformations and von Mises stress prediction (color, yellow is high) at $t=100$, $t=200$, and $t=400$, respectively. In each figure, rows correspond to truncation steps $k_{\text{tr}}=3,5,10,20$. Using a low truncation step $k_{\text{tr}}$ increases inference speed and enables robust predictions of the global deformation modes. However, it also causes local mesh degradation due to the accumulation of high-frequency errors, as observed in MSE trained one step models [9].
(a) ImpactPlate.
(b) DeformingPlate.
Figure 15: Comparison of the rollout displacement RMSE and displacement gradient RMSE for different ROBIN setting (denoising stride $m$ / truncation step $k_{\text{tr}}$) on a) ImpactPlate and b) DeformingPlate. Early diffusion steps substantially reduce the displacement RMSE and later steps the local displacement gradient RMSE. The fast inference setting (1/20) remains accurate and yields slightly lower displacement RMSE on ImpactPlate, but also slightly higher gradient RMSE.
Quantitative results.
Table˜3 lists the quantitative results of all experiments considered in this work and used for the visualizations.
Table 3: Quantitative results on BendingBeam, ImpactPlate, DeformingPlate, and the large-mesh dataset BendingBeamLarge. We report displacement RMSE $10^{-3}$], gradient RMSE of the displacement field $10^{-3}$], and rollout wall-clock time [s]. Values are mean $\pm$ std over the test set across 5 seeds (solver timings excluded). The row Numerical solving gives the average runtime of the high-fidelity solver (maximum in parentheses). Baselines contrasts different models, Variants sweep the ROBI settings (denoising stride / truncation step) and additionally report Gradient RMSE, Ablations remove ROBIN components, and Generalization to large meshes evaluates the upscaling capabilities of ROBIN.
BendingBeam ImpactPlate DeformingPlate
Numerical solving Time [s] Time [s] Time [s]
46.2 (max. 186.1) 742.6 [22] 1157.2 [5]
Baselines
RMSE $10^{-3}$] Time [s] RMSE $10^{-3}$] Time [s] RMSE $10^{-3}$] Time [s]
ROBIN 29.00 $\pm$ 1.00 15.03 $\pm$ 0.04 12.33 $\pm$ 0.96 4.94 $\pm$ 0.01 4.98 $\pm$ 0.33 61.60 $\pm$ 0.29
HCMT 121.53 $\pm$ 1.87 25.82 $\pm$ 0.28 19.57 $\pm$ 0.38 4.33 $\pm$ 0.05 8.04 $\pm$ 0.13 31.57 $\pm$ 0.14
MGN 189.52 $\pm$ 70.87 21.40 $\pm$ 0.15 54.07 $\pm$ 5.88 2.93 $\pm$ 0.03 8.76 $\pm$ 0.29 24.42 $\pm$ 0.22
BSMS 141.98 $\pm$ 7.96 4.70 $\pm$ 0.04 63.52 $\pm$ 32.36 2.16 $\pm$ 0.01 13.59 $\pm$ 5.21 13.28 $\pm$ 0.07
Variants (denoising stride / truncation step)
RMSE $10^{-3}$] Time [s] RMSE $10^{-3}$] Time [s] RMSE $10^{-3}$] Time [s]
20/20 28.90 $\pm$ 1.91 162.43 $\pm$ 0.89 14.43 $\pm$ 1.93 22.61 $\pm$ 0.36 4.96 $\pm$ 0.36 190.10 $\pm$ 1.33
5/20 28.99 $\pm$ 1.93 43.39 $\pm$ 0.52 13.87 $\pm$ 1.90 7.83 $\pm$ 0.06 4.92 $\pm$ 0.35 83.61 $\pm$ 0.16
1/20 29.00 $\pm$ 1.00 15.03 $\pm$ 0.04 12.33 $\pm$ 0.96 4.94 $\pm$ 0.01 4.98 $\pm$ 0.33 61.60 $\pm$ 0.29
1/10 30.35 $\pm$ 2.41 11.03 $\pm$ 0.10 12.38 $\pm$ 0.92 2.90 $\pm$ 0.01 5.18 $\pm$ 0.35 33.71 $\pm$ 0.10
1/5 35.81 $\pm$ 3.30 9.88 $\pm$ 0.19 13.01 $\pm$ 0.98 1.91 $\pm$ 0.01 5.73 $\pm$ 0.30 20.24 $\pm$ 0.04
1/3 42.83 $\pm$ 4.54 9.57 $\pm$ 0.11 14.58 $\pm$ 1.40 1.55 $\pm$ 0.01 6.42 $\pm$ 0.23 14.85 $\pm$ 0.11
1/2 48.94 $\pm$ 7.25 9.47 $\pm$ 0.11 15.79 $\pm$ 1.71 1.42 $\pm$ 0.00 7.26 $\pm$ 0.18 12.59 $\pm$ 0.13
1/1 64.23 $\pm$ 11.86 9.49 $\pm$ 0.12 21.46 $\pm$ 2.12 1.32 $\pm$ 0.01 10.79 $\pm$ 0.52 7.11 $\pm$ 0.04
Gradient RMSE $10^{-3}$] Gradient RMSE $10^{-3}$] Gradient RMSE $10^{-3}$]
20/20 12.75 $\pm$ 0.70 1.92 $\pm$ 0.07 28.98 $\pm$ 1.56
5/20 12.74 $\pm$ 0.65 1.89 $\pm$ 0.08 28.83 $\pm$ 1.57
1/20 12.77 $\pm$ 0.72 2.55 $\pm$ 0.33 29.10 $\pm$ 1.50
1/10 37.58 $\pm$ 4.71 2.61 $\pm$ 0.30 41.59 $\pm$ 1.76
1/5 94.00 $\pm$ 20.90 3.68 $\pm$ 0.37 73.13 $\pm$ 5.64
1/3 147.97 $\pm$ 65.13 6.07 $\pm$ 0.56 110.96 $\pm$ 22.33
1/2 200.44 $\pm$ 128.29 8.61 $\pm$ 1.02 150.33 $\pm$ 48.48
1/1 324.05 $\pm$ 285.65 23.41 $\pm$ 4.85 157.30 $\pm$ 8.34
Ablations
RMSE $10^{-3}$] RMSE $10^{-3}$] RMSE $10^{-3}$]
ROBIN 29.00 $\pm$ 1.00 12.33 $\pm$ 0.96 4.98 $\pm$ 0.33
10 diff. steps 33.28 $\pm$ 2.43 14.58 $\pm$ 1.99 5.86 $\pm$ 0.61
5 diff. steps 41.33 $\pm$ 6.27 15.60 $\pm$ 1.74 6.40 $\pm$ 0.56
w/o shared layer 66.05 $\pm$ 7.36 12.88 $\pm$ 1.58 5.65 $\pm$ 0.26
State prediction 596.38 $\pm$ 70.93 130.50 $\pm$ 30.80 137.61 $\pm$ 30.54
w/o hierarchy 122.38 $\pm$ 1.52 57.02 $\pm$ 3.48 12.04 $\pm$ 0.24
w/o diffusion 43.83 $\pm$ 1.16 99.47 $\pm$ 47.30 9.78 $\pm$ 2.12
HCMT model 111.05 $\pm$ 1.15 29.18 $\pm$ 1.43 14.86 $\pm$ 0.42
BendingBeamLarge
Numerical solving Time [s]
108.30 (max. 4248.01)
Generalization to large meshes
RMSE $10^{-3}$] Time [s]
Pre-trained 77.59 $\pm$ 4.97 30.94 $\pm$ 0.63
From scratch 215.50 $\pm$ 12.14 30.95 $\pm$ 0.58